We have the following information about ATT's bonds.

Semi-annual

coupon rate=0

Par value=$1000

Market price=$460

Time-to-maturity=15 years

What is the YTM on ATT's bonds?

Group of answer choices

5.24%

4.31%

7.22%

1.94%
Answer #1

Using financial calculator

N=15*2=30

PMT=0

FV=1000

PV=-460

CPT I/Y=2.62%

Hence, yield to maturity=2.62%*2=5.24%